# Seed-Music: A Unified Framework for High Quality and Controlled Music Generation

### v2

@@ 2 Literature Review
Each approach presents unique advantages and limitations depending on the task. Active research continues to focus on identifying optimal representations for music audio, including compression methods that balance short-term and long-term sequences, encoding techniques that can jointly capture melodic, harmonic, and rhythmic information, and strategies for navigating the trade-offs between high-level semantic and low-level acoustic features.
@@ 3 Method
Representation Compression Interpretability Renderer friendly Generator friendly
Audio token yes no maybe yes
Symbolic music yes yes no yes
Vocoder latent maybe no yes maybe
@@ 3.2 Symbolic Token-based Pipeline
Prior efforts have proposed algorithms for melody generation (Ju et al., 2021; Zhang et al., 2022). However, they lack explicit phoneme- and note-aligned information crucial for vocal music generation. Moreover, they remain limited to symbolic music generation without the capability for audio rendering. In a different line of research, there are task-specific prior works studying the approaches to steer music audio generation through musically interpretable conditions like harmony (Copet et al., 2024), dynamics, and rhythm (Wu et al., 2023). Inspired by how jazz musicians use lead sheets to outline a composition’s melody, harmony and structure, we introduce “lead sheet tokens” as a symbolic music representation. We highlight the key components, benefits, and limitations of lead sheet tokens compared to audio tokens below.
@@
  • As the lead sheet tokens lack acoustic feature characterization, we scale up the token-to-latent diffusion model in the cascaded Renderer to achieve the same end-to-end performance as the audio token-based system.
@@ 4.1 Lyrics2Song
In both modes, our model demonstrates a strong ability to maintain coherence between the input lyrics and the inherent lyrics in the audio reference, even without the aid of automatic lyrics transcription. When the input lyrics are structurally and semantically similar to those in the audio reference, the model tends to resample the melody and structure from the reference. However, when the input lyrics differ significantly in style (e.g., language, structure, rhyme), the coherence weakens. Despite this, the model effectively maintains natural rhythm patterns, instrumentation, vocal quality, and overall musical motifs.

### v3

@@ 2 Literature Review
Each approach offers distinct advantages and limitations depending on the task. Ongoing research continues to explore optimal representations for music audio, focusing on aspects such as compression methods that balance between short-term features and long-term coherence, disentanglement mechanisms for melodic, harmonic, and rhythmic characteristics, and the trade-offs between high-level semantic and low-level acoustic features.
@@ 3 Method
Representation Compression Interpretability Generator-Friendly Renderer-Friendly
Audio token yes no yes maybe
Symbolic music yes yes yes no
Vocoder latent maybe no maybe yes
@@ 3.2 Symbolic Token-based Pipeline
Prior efforts have proposed algorithms for melody generation (Ju et al., 2021; Zhang et al., 2022). However, they lack explicit phoneme- and note-aligned information crucial for vocal music generation. Moreover, they remain limited to symbolic music generation without the capability for audio rendering. In a different line of research, there are task-specific prior works studying the approaches to steer music audio generation through musically interpretable conditions like harmony (Copet et al., 2024), dynamics, and rhythm (Wu et al., 2023). Inspired by how jazz musicians use lead sheets to outline a composition’s melody, harmony and structure, we introduce “lead sheet tokens” as the symbolic music representation. We highlight the key components, benefits, and limitations of lead sheet tokens compared to audio tokens as follows.
@@
  • As the lead sheet tokens lack acoustic feature characterization, we need to scale up the token-to-latent diffusion model in the cascaded Renderer to achieve the same end-to-end performance as the audio token-based system.
@@ 4.1 Lyrics2Song
In both modes, our model demonstrates a strong ability to maintain coherence between the input lyrics and the inherent lyrics in the audio reference, even without the aid of automatic lyrics transcription. When the input lyrics are structurally and semantically similar to those in the audio reference, the model tends to resemble the melody and structure from the reference. However, when the input lyrics differ significantly in style (e.g., language, structure, rhyme), the coherence weakens. Despite this, the model effectively maintains natural rhythm patterns, instrumentation, vocal quality, and overall musical motifs.